myininaya (myininaya):

This question is from a James Stewart Calculus book: Suppose f is a function that satisfies the equation f(x+y)=f(x)+f(y)+x^2y+xy^2 for all real numbers x,y. Also suppose as x->0 , f(x)/x->1. Then we have 3 questions: a,b, and c. a. Find f(0). b. Find f'(0). c. Find f'(x).

OpenStudy (perl):

nice question, hmmm

OpenStudy (anonymous):

ooh ooh i got the first one!

myininaya (myininaya):

oops let me trying posting that one more time

OpenStudy (anonymous):

i got \(f(0)=0\) right?

myininaya (myininaya):

right

OpenStudy (ksaimouli):

(d) deals with partial derivatives?

OpenStudy (ksaimouli):

Never mind we can use l'hopitals rule

myininaya (myininaya):

This is all the problem gave. But I can tell you partial derivatives have not been covered before where this question is placed in the book.

myininaya (myininaya):

I'm actually having trouble determine if f'(0) exists or not. Like I found a contradiction in something I was doing.

myininaya (myininaya):

I know f'(0)=1 or f'(0) dne.

myininaya (myininaya):

I assumed f'(0)=1 but then I got a contradiction which led me to believe f'(0) dne

myininaya (myininaya):

But I can't post my latex because I can't find a problem in what I typed yet

myininaya (myininaya):

\[ \text{ We have } f(x+y)=f(x)+f(y)+x^2y+xy^2 \\ \text{ Let } y=0 \text{ so } f(x)=f(x)+f(0) \text{ => } f(0)=0 \\ \text{ Since } \lim_{x \rightarrow 0} \frac{f(x)}{x}=1 \] \[ \text{ and we have} f(0)=0 \text{ then we also have } \\ \lim_{x \rightarrow 0 } f’(x)=1 \\ \text{ From this, I can only tell either we have } f’(0)=1 \text{ or } \\ f’(0) \text{ does not exist} \\ \] \[ \text{ I also know } f \text{ is an odd function: } \\ \text{ let } y=-x \\ f(x+(-x))=f(x)+f(-x)+x^2(-x)+x(-x)^2 \\ f(0)=f(x)+f(-x) \\ 0=f(x)+f(-x)\\ -f(x)=f(-x)\\ \] \[\text{ But I don’t think this helps me to determine f’(0) } \\ \text{ I think } f \text{ is continuous because it said that equation was for all real numbers}\\ u=x+y \\ \text{ but I’m not sure } f \text{ is differentiable for all real numbers } u \\ \text{ We also know from } f \text{ being odd that } \\ f’ \text{ will be even } \\ \] \[ \text{ So I’m going to differentiate } f \text{ now } \\ f’(x+y)=f’(x)+f’(y)+x^2+y^2+4xy \\ \text{ Let } y=-x \text{ so we have assuming that } f’\text{ exists at } 0 \\ f’(x+(-x))= f’(x)+f’(-x)+x^2+x^2-4x^2=2f’(x)-2x^2 \\ \text{ so we have } f’(0)=2f’(x)-2x^2 \\ \text{ so } f’(x)=\frac{ f’(0)+2x^2 }{2}\\ \] \[ \text{ which means going back to } \\ \lim_{x \rightarrow 0 } f’(x)=1 \\ \text{ we have } \\ \lim_{x \rightarrow 0 } \frac{ f’(0)+2x^2 }{2}=1 \\ \lim_{x \rightarrow 0 } (f’(0)+2x^2)=2 \\ \text{ but if } f’(0)=1 \text{ then we have a contradiction because} \\ 1+0 \neq 2 \\ \text{ So this leads me to think } f’(0) \text{ does not exist } \] \]

myininaya (myininaya):

I did not say derivative of y is y' because I think it doesn't depend on any number like I'm saying y is not a function of x

myininaya (myininaya):

or I mean on any variable (number ) whatever

myininaya (myininaya):

wait ... f'(0) doesn't have to be 1 or does not exist it could be some other number

myininaya (myininaya):

if I wanted f'(0) to exist it would have to be 2

myininaya (myininaya):

in order for that one limit to hold

myininaya (myininaya):

\[\lim_{x \rightarrow 0}(f'(0)+2x^2) =2 \] since x->0 then 2x^2->0 so f'(0) is 2.. so f'(x)=1+x^2

myininaya (myininaya):

Let me know if you guys disagree on anything

ganeshie8 (ganeshie8):

im still at half way of the work, i like the odd function part xD

myininaya (myininaya):

i do too

myininaya (myininaya):

I thought it was the cutest thing about the problem

myininaya (myininaya):

lol but if f'(x)=1+x^2 then f'(0)=1+0=1 not 2 so f'(0) does not exist?

OpenStudy (xapproachesinfinity):

this is really good

myininaya (myininaya):

well my answer is not really good

myininaya (myininaya):

but i think the problem is good

myininaya (myininaya):

The weird thing is if you have completed cal 1 all the way up passed implicit differentiation you should be able to do this problem since it is right after the chapter 3 review.

OpenStudy (xapproachesinfinity):

So far what you did is really nice^_^

OpenStudy (matimaticas):

I think you guys get buzzed off of math lol

ganeshie8 (ganeshie8):

myininaya i am struggling with understanding below derivative - \[\text{ So I’m going to differentiate } f \text{ now } \\ \color{Red}{f’(x+y)=f’(x)+f’(y)+x^2+y^2+4xy }\\ \text{ Let } y=-x \text{ so we have assuming that } f’\text{ exists at } 0 \\ f’(x+(-x))= f’(x)+f’(-x)+x^2+x^2-4x^2=2f’(x)-2x^2 \\ \text{ so we have } f’(0)=2f’(x)-2x^2 \\ \text{ so } f’(x)=\frac{ f’(0)+2x^2 }{2}\\\] how do we differentiate when with respect to (x+y) ?

myininaya (myininaya):

I was feeling icky about that part.

myininaya (myininaya):

I wasn't sure about it. I was looking at if as if f depended on x and y Like f(x)=x^2 we could say this f'(x)=2x or like if w ehad f(y)=y^2 we could say f'(y)=2y

myininaya (myininaya):

If that makes sense

myininaya (myininaya):

But I do know we have f(x+y) and if f(x+y)=x+y then f'(x+y)=1

myininaya (myininaya):

so yeah I think want to scratch that

ganeshie8 (ganeshie8):

I thought you have deduced that part using the fact that f'(x) is even, but still trying to look for strong evidence x^2y + xy^2 = xy(x+y) we need to use partials or some other means to break it down i think

myininaya (myininaya):

This question comes way before partials though

myininaya (myininaya):

ok I think I got it...

myininaya (myininaya):

maybe

myininaya (myininaya):

\[f(\frac{x}{2}+\frac{x}{2})=f(\frac{x}{2})+f(\frac{x}{2})+(\frac{x}{2})^2\frac{x}{2}+\frac{x}{2}(\frac{x}{2})^2\]

myininaya (myininaya):

not we can differentiate with respect to one variable

myininaya (myininaya):

so we have \[f'(x)=f'(\frac{x}{2})+\frac{3x^2}{2}\]

ganeshie8 (ganeshie8):

Oh nice, but we can't plugin x=0 :/

myininaya (myininaya):

yeah so sad

myininaya (myininaya):

i'm thinking maybe we could use the f(x)/x->1 as x->0 more ... maybe \[f'(0)=\lim_{x \rightarrow 0}\frac{f(x)-f(0)}{x-0}=1 \] so f'(0) is 1

myininaya (myininaya):

since f(0) is 0

myininaya (myininaya):

so now we need to find f'(x)

ganeshie8 (ganeshie8):

Wow! that looks like a clever idea

ganeshie8 (ganeshie8):

we can use all our previous discoveries here i feel

myininaya (myininaya):

something about f being off and f' being even?

ganeshie8 (ganeshie8):

\[ \large f'(x)=\lim_{x \rightarrow y}\frac{f(x)-f(y)}{x-y }\]

ganeshie8 (ganeshie8):

or maybe let x->-y so that we can use f(-y) = -f(y) to our advantage : \[ \large f'(x)=\lim_{x \rightarrow -y}\frac{f(x)-f(-y)}{x-(-y) }\] \[ \large f'(x)=\lim_{x \rightarrow -y}\frac{f(x)+f(y)}{x+y }\]

ganeshie8 (ganeshie8):

\[ \large f'(x)=\lim_{x \rightarrow -y}\frac{f(x+y ) - xy(x+y)}{x+y }\] \[ \large f'(x)=\lim_{x \rightarrow -y}\frac{f(x+y )}{x+y } - xy \] \[ \large f'(x)=1 +y^2\]

ganeshie8 (ganeshie8):

how did we endup f'(x) in terms of y hmm

myininaya (myininaya):

well both of your limits tend to y

myininaya (myininaya):

I think it might be f'(y)=1+y^2 so f'(x)=1+x^2 which is what I actually had earlier doing things in an illegal way

ganeshie8 (ganeshie8):

I see, need to check this with fresh mind in the morning again. thanks for the beautiful question myininaya :)

myininaya (myininaya):

Thanks for looking at @ganeshie8

myininaya (myininaya):

I like looking at the problem plus section at the end of each section in the cal 1-3 book